OpenStudy (anonymous):

Jack invested some money in a bank at a fixed rate of interest compounded annually. The equation below shows the value of his investment after x years f(x) = 300(1.02)^x What was the average rate of change of the value of Jack's investment from the third year to the fifth year? WILL FAN FOR THE ANSWER PLZ HELP

OpenStudy (naga447):

It's one of these can you figure out which one it is? 6.43 dollars per year 8.24 dollars per year 12.86 dollars per year 14.26 dollars per year

OpenStudy (anonymous):

Do you guys know the answer to this question?

OpenStudy (wmj259):

Use this total = principal * (1 + rate) ^time (notice we need exponentiation here.) 300=300*(1+.02)^(0) 306=300*(1+.02)^(1) 312.12=300*(1+.02)^(2) 318.36=300*(1+.02)^(3) 324.73=300*(1+.02)^(4) 331.22=300*(1+.02)^(5) Year 5 - Year 3 = 331.22-318.36=12.86